Question:If in stopping a car of mass 800 kg by applying brakes,4.18 kcal of heat is produced,then what was the speed of the car before the applying the brakes ?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
(mv^2)/2 = 4.18 × 10^3 × 4.184.
v^2 = 17489.12*2 / 800 = 21.8614 *2
Therefore,
v = 6.612 m/s^2.
 
 
Regards
Arun (askIITians forum expert)